OpenStudy (anonymous):

Hexagon A is a regular hexagon. The total length of all the sides of the hexagon is 24 inches. Hexagon A is dilated about its center to create Hexagon B. The length of each side of Hexagon B is 4 4/5 inches. By what factor was Hexagon A dilated to create Hexagon B? A. 5/6 B. 1 1/2 C. 2 D. 4/5

OpenStudy (luigi0210):

Welcome to Openstudy!

OpenStudy (anonymous):

Glad to be here? :D

OpenStudy (luigi0210):

Okay, so do you know how many sides a hexagon has?

OpenStudy (anonymous):

6

OpenStudy (luigi0210):

Correct, so for the first Hexagon, hexagon A, it has a total of 24 inches. To find the side length of each, we need to divide that by 6. So what's: \[\frac{24}{6}\]

OpenStudy (anonymous):

4

OpenStudy (luigi0210):

Right, so each side of hexagon A is 4 inches. Hexagon B has sides of 4.5 inches. What's the difference there?

OpenStudy (anonymous):

4/5 s? (Hexagon B is 4 4/5, actually) c:

OpenStudy (luigi0210):

Oh, right sorry I mis-read that. Thanks for the correction. So what's the difference between \[4\frac{4}{5}-4=?\]

OpenStudy (anonymous):

4/5 c:

OpenStudy (luigi0210):

Yup, so unless my math skills have suddenly vanished into thin air, that should be the answer.

OpenStudy (austinl):

Actually wait. \(\dfrac{24}{6}=l_A=4=\dfrac{20}{5}\\ l_b=\dfrac{24}{5}\\ \text{factor of dilation}=\dfrac{l_A}{l_B}=\dfrac{\dfrac{20}{5}}{\dfrac{24}{5}}=\dfrac{5}{6}\)

OpenStudy (luigi0210):

Thanks Austin, I probably should of asked what dilation was.. ._.